OpenStudy (anonymous):

I need help on math homework, I'm not asking you to do it for me just help. Math is my worst subject but it says to choose the equation of the line passing through the point (1,-5) and perpendicular to y=1/5 x -10. A. y= 5x + 10 B. y=5x C. Y= -5x -10 D. y= -5x

jimthompson5910 (jim_thompson5910):

The slope of y=1/5 x -10 is what?

jimthompson5910 (jim_thompson5910):

If you're not sure, then remember that y = mx+ b is the general form where m is the slope ex: y = 3x + 5 has a slope of m = 3 ex: y = 1/9x - 10 has a slope of m = 1/9

OpenStudy (anonymous):

i love you kesha

OpenStudy (anonymous):

I'm so confused

OpenStudy (anonymous):

i love u

jimthompson5910 (jim_thompson5910):

what's the number in front of the x variable

jimthompson5910 (jim_thompson5910):

in the equation y=1/5 x -10

OpenStudy (anonymous):

1/5 right?

jimthompson5910 (jim_thompson5910):

so the slope of y=1/5 x -10 is 1/5

jimthompson5910 (jim_thompson5910):

to find the perpendicular slope, you flip the fraction to get 5/1 = 5 then you flip the sign to get -5

jimthompson5910 (jim_thompson5910):

so the perpendicular slope is -5

jimthompson5910 (jim_thompson5910):

this means your answer will have a slope of -5

OpenStudy (anonymous):

im attracted to you

jimthompson5910 (jim_thompson5910):

so y = mx+b y = -5x + b ... plug in the perpendicular slope -5 = -5(1) + b ... plug in x = 1 and y = -5 -5 = -5 + b solve for b and tell me what you get

OpenStudy (anonymous):

So the answer is B?

OpenStudy (anonymous):

yes it is

jimthompson5910 (jim_thompson5910):

close, but notice that the slope is negative

jimthompson5910 (jim_thompson5910):

no the answer isn't B

OpenStudy (anonymous):

jk i knew that. i love you

jimthompson5910 (jim_thompson5910):

ok you need to either help or get out bobby, this is for math only save that for the chat or private messaging

jimthompson5910 (jim_thompson5910):

or better yet, do it on facebook

OpenStudy (anonymous):

okkk screw me then

OpenStudy (anonymous):

its D right?

jimthompson5910 (jim_thompson5910):

yes it's y = -5x

jimthompson5910 (jim_thompson5910):

-5 = -5 + b -5+5 = -5 + b + 5 0 = b b = 0 so m = -5 b = 0 ---------- y = mx + b y = -5x + 0 y = -5x

OpenStudy (anonymous):

omg thank you so much! <3 i kinda understand it

jimthompson5910 (jim_thompson5910):

ok just keep practicing it and you'll get much better at it